OpenStudy (anonymous):

What is the rational number equivalent to 1.28? The 1.28 has a line above the 28

OpenStudy (phi):

the line over the 28 means the 28 repeats forever. 1.282828.... and so on let x be the rational number 1.28... we can use this trick: 100*1.282828....= 128.282828... (the decimal 28 part repeats) 100x = 128.28... next: 100x - x = 128.282828... - 1.282828... the .282828... part will be subtracted away 99x = 127 divide both sides by 99 to get x= 127/99